The current I in an electrical conductor varies inversely as the resistnace R of the conductor. The current is 7 amperes when the resistance is 990 ohms. What is the current when the resistance is 510 ohms?
Answer by Alan3354(69443)

The current I in an electrical conductor varies inversely as the resistnace R of the conductor. The current is 7 amperes when the resistance is 990 ohms. What is the current when the resistance is 510 ohms?
------------
I = k/R
I*R = k (constant)
------------
7*990 = I*510